I am working on a new lesson and I an need of some help. I need to write an equation of the line that is parallel to the given line and passes through the given point.
Y=2/3X-2, (2,1)
Please help!

Answer by stanbon(75887)

write an equation of the line that is parallel to the given line and passes through the given point.
Y=2/3X-2, (2,1)
--------------
The given line has a slope of 2/3.
Any line parallel to it must have a slope of 2/3.
----
The form of the equation you want is y = mx + b
---
You know m = 2/3 and you know y = 1 when x = 2:
So, solve for "b":
----
1 = (2/3)*2 + b
1 = (4/3) + b
b = -1/3
---------------------
Equation:
y = (2/3)x - (1/3)
==========================
Cheers,
Stan H.